3. Solve by substitution or elimination method:
-2x + 6y = 19
10x – 30y = -15

Answer by J2R2R(94)

Solve by substitution or elimination method:
-2x + 6y = 19
10x - 30y = -15

This looks like it is either a trick question or not correctly written.

Multiply the first equation by -5

10x - 30y = -95

This contradicts the second equation.

Sorry but it cannot be solved.